OpenStudy (anonymous):

Indicate in standard form the equation of the line through the given points. K(6, 4), L(-6, 4)

jimthompson5910 (jim_thompson5910):

same idea as last time....but...we don't have the slope

jimthompson5910 (jim_thompson5910):

so you need to find the slope first

OpenStudy (anonymous):

yeah very similar

jimthompson5910 (jim_thompson5910):

how do you find the slope

OpenStudy (anonymous):

not sure :(

jimthompson5910 (jim_thompson5910):

use the slope formula m = (y2 - y1)/(x2 - x1)

jimthompson5910 (jim_thompson5910):

(x1,y1) is your first point (x2,y2) is your second point

OpenStudy (anonymous):

0

OpenStudy (anonymous):

/-12

jimthompson5910 (jim_thompson5910):

so the slope is?

OpenStudy (anonymous):

-12

jimthompson5910 (jim_thompson5910):

what did you get for the numerator

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

as long as the denominator is not 0, the slope will always be 0 if the numerator is 0 since 0/x = 0 x is any number but 0

jimthompson5910 (jim_thompson5910):

so the slope is 0 instead of -12

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

now you have enough info to find the equation

OpenStudy (anonymous):

hmmmm i'm lost

OpenStudy (anonymous):

don't know where to start

jimthompson5910 (jim_thompson5910):

remember you use y-y1 = m(x-x1)

OpenStudy (anonymous):

oh right!

jimthompson5910 (jim_thompson5910):

that's why you needed the slope first, you weren't given m (like last time) so you needed to find m before you could move on

OpenStudy (anonymous):

hmmm so would it be y-4=0?

jimthompson5910 (jim_thompson5910):

close, they want it in standard form

jimthompson5910 (jim_thompson5910):

standard form has the constant all by itself on the right side

OpenStudy (anonymous):

-y+4?

jimthompson5910 (jim_thompson5910):

y-4 = 0 was close, you just need to get it into the form Ax+By = C

OpenStudy (anonymous):

y=4?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

that's it

OpenStudy (anonymous):

Great! thanks a lot you've been a big help

jimthompson5910 (jim_thompson5910):

glad I could be